OpenStudy (yas123):

write an equation in slope intercept form for the line containing the indicated points. (2,4) and (3,5) help !

OpenStudy (anonymous):

i just solved one just like this for you..

OpenStudy (anonymous):

(y2-y1)/(x2-x1) (5-4)/(3-2) 1/1 Slope = 1 y = x + b 4 = 2 + b b = 2 y = x + 2

OpenStudy (anonymous):

y=mx+b m=slope y=1(x) + b is how I look at it but Pk51 looks right

OpenStudy (yas123):

fishcx : i know but i keep getting diffrent answers